OpenStudy (anonymous):

state and explain with derivation ,the balanced condition of wheatstone bridge.

OpenStudy (aravindg):

i think its r1/r2=r3/r4

OpenStudy (aravindg):

its there in ncert text well explained

OpenStudy (anonymous):

i did'nt got it well :(

OpenStudy (aravindg):

:)

OpenStudy (aravindg):

did u search the net ?? when i prepared for KVPY i got a vry nic article on this ,but i dont hav that nw

OpenStudy (anonymous):

not a link?

OpenStudy (anonymous):

The balance i9n the wheatstone bridge is attained when the current through the galvanometer reads zero. It may be given as P/Q=R/S